OpenStudy (anonymous):

]Identify the correct x intercept and y intercept of the equation x + 7y = -21.

OpenStudy (anonymous):

to find x-intercepts, set y = 0 when y = 0 x + 7*0 = -21 x = -21 when x=0 0 + 7y = -21 y = -21/7 = -3

OpenStudy (anonymous):

to find y-intercepts, set x=0. that's what I did above.